# BLIPs: Bayesian Learned Interatomic Potentials

### v1

@@ 1 Introduction
Deep learning is transforming computational chemistry and materials science, enabling advances in molecular design (Özçelik et al.,, 2024; Eijkelboom et al.,, 2024; Zeni et al.,, 2023; Merchant et al.,, 2023), simulation-based modelling (Batatia et al.,, 2022; Wood et al.,, 2025; Neumann et al.,, 2024), and particle-based systems (Kipf et al.,, 2018; Alkin et al.,, 2024; Zhdanov et al.,, 2025). At the forefront of this progress are Message Passing Neural Networks (MPNNs) (Scarselli et al.,, 2008; Kipf and Welling,, 2016; Gilmer et al.,, 2017), which have emerged as core architectures for learning in atomistic and molecular domains. Their ability to natively incorporate essential physical symmetries, such as translational, rotational, and permutation invariance or equivariance (Cohen and Welling,, 2016; Weiler et al.,, 2023; Bronstein et al.,, 2017; Cohen et al.,, 2019), makes them particularly well-suited for modelling the complex interactions that define atomic-scale systems. Consequently, MPNNs are now widely adopted in the development of Machine Learning Interatomic Potentials (MLIPs), where they serve as efficient and accurate surrogates for quantum and classical simulations. These models enable large-scale atomistic simulations with near ab initio accuracy, dramatically accelerating research in materials discovery, catalysis, and molecular dynamics (Wood et al.,, 2025; Neumann et al.,, 2024; Batatia et al.,, 2022).
@@
However, like most deep learning models, MLIPs are vulnerable to failure, particularly when faced with out-of-distribution (OOD) data. This issue is further worsened in scientific domains, where generating labelled data often requires expensive and time-consuming simulations (Papamarkou et al.,, 2024). Indeed, for simulation-based chemistry, training data is often obtained from Density Functional Theory (DFT) calculations, which scale cubically with system size, making it impractical to generate data for larger molecules or materials. Moreover, the chemical space is vast and inherently undersampled, making distributional shifts inevitable (Özçelik et al.,, 2025; Tan et al.,, 2023). These challenges make principled Uncertainty Quantification (UQ) essential for enabling error-aware simulations or guiding active learning pipelines. Among existing approaches, Deep Ensembles (Lakshminarayanan et al.,, 2017) are commonly used for UQ in MLIPs and have shown strong empirical performance compared to other UQ methods (Tan et al.,, 2023). However, they require training and storing multiple models, making them computationally and memory-intensive, particularly in large-scale or high-throughput scientific workflows. Other simpler UQ methods, such as MC Dropout (Gal and Ghahramani,, 2016) or mixture models, tend to be less memory-intensive but provide worse uncertainty estimates, and their predictive accuracy may degrade compared to standard deterministic models.
To overcome these limitations, we introduce Bayesian Learned Interatomic Potentials (BLIPs), a general variational Bayesian method that is scalable, accurate, and provides principled, well-calibrated uncertainty estimates for MLIPs. BLIP can be applied both when training or fine-tuning pretrained machine learning interatomic potentials. It builds on an adaptive version of Variational Dropout (Kingma et al.,, 2015), where the MPNNs’ weights are stochastically perturbed depending on the network’s input data. BLIP does not depend on the chosen MPNN architecture and seamlessly integrates with a vast family of invariant and equivariant models. Importantly, our approach adds minimal computational overhead compared to deterministic models in both training and predictive inference, while providing well-calibrated uncertainties and improvement in predictive accuracy. We evaluate our approach on various simulation-based chemistry tasks, focusing on small and large-scale systems. In both domains, we show that training with BLIP not only improves predictive performance but also produces meaningful uncertainty estimates. Furthermore, we demonstrate that fine-tuning a single pretrained MLIP with BLIP not only consistently improves predictive performance but also outperforms an ensemble of fine-tuned models, while simultaneously providing reliable uncertainty estimates. The code accompanying this work is available at https://github.com/dario-coscia/blip.
@@ 2.1 Message Passing Neural Networks
Message Passing Neural Networks (Scarselli et al.,, 2008; Kipf and Welling,, 2016; Gilmer et al.,, 2017) are a class of (permutation equivariant) neural networks that process graph-structured data. Given a Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $v_{i}\in\mathcal{V}$, and edges $e_{ij}\in\mathcal{E}$, the MPNNs operations for each layer $l=0,\dots,L-1$ can be summarised as:
@@
where $\bm{h}_{i}^{l}$ is the feature vector for node $i$ at layer $l$, $a_{ij}$ are the edge attributes, and $\bigoplus_{j\in\mathcal{N}(i)}$ is a permutation invariant pooling operation over the node $i$ neighbors $\mathcal{N}(i)$ (e.g. sum or mean). Finally, $M_{l},U_{l}$ are learnable message and update functions usually approximated with small neural networks.
@@ 2.2 Bayesian Models and Uncertainty Quantification
Bayesian models have a long history in Deep Learning (Hinton and Van Camp,, 1993; Welling and Teh,, 2011; Graves,, 2011; Blundell et al.,, 2015; Gal and Ghahramani,, 2016), providing a principled framework for modelling uncertainty through probabilistic inference. Instead of treating neural network weights as fixed parameters, Bayesian approaches represent them as probability distributions, $\bm{\omega}\sim p(\bm{\omega})$. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ of $N$ i.i.d. input-output observations, the goal is to infer the posterior distribution over weights, $p(\bm{\omega}\mid\mathcal{D})$, using Bayes’ theorem (Bayes,, 1763). Predictive inference is then performed via Bayesian model averaging, where the predictive distribution is given by $p(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{p(\bm{\omega}\mid\mathcal{D})}[p(\mathbf{y}\mid\mathbf{x},\bm{\omega})]$. However, due to the high dimensionality of weight spaces in modern deep networks, exact Bayesian inference is computationally intractable, necessitating approximate methods. Among these, Deep Ensembles approximate the posterior by training multiple networks with different random initialisations and aggregating their predictions (Lakshminarayanan et al.,, 2017) and has been shown to provide state-of-the-art results in many applications (Fort et al.,, 2019; Tan et al.,, 2023). Other widely used approaches include Variational Inference (Blundell et al.,, 2015; Kingma et al.,, 2015), which approximates the posterior with a tractable parametric distribution optimised via the evidence lower bound (ELBO), and Monte Carlo Dropout, which interprets dropout at test time as Bayesian inference (Gal and Ghahramani,, 2016).
@@ 2.3 Machine Learning Interatomic Potentials
Density Functional Theory (DFT) (Hohenberg and Kohn,, 1964; Kohn and Sham,, 1965; Parr,, 1989) is a widely used quantum mechanical method for calculating a range of properties for various atomic systems. Under the Born-Oppenheimer approximation (Born and Heisenberg,, 1985) of DFT, the potential energy surface (PES) $E$ of an atomic system can be computed as a function of positions $\bm{r}$ and atomic numbers $\bm{Z}$. Forces can be obtained by taking the negative gradient with respect to atom-positions, $\bm{F}=-\nabla_{\bm{r}}E(\bm{r},\bm{Z})$. DFT calculations are computationally expensive, due to their cubic scaling with respect to the number of electrons, and to overcome this limitation, Machine Learning Interatomic Potentials (MLIPs) have emerged as efficient surrogates for it, aiming to approximate the PES accurately and their corresponding forces, using machine learning models. Among the broad family of machine learning models, (equivariant) MPNNs (Batatia et al.,, 2022; Fu et al.,, 2025; Neumann et al.,, 2024; Rhodes et al.,, 2025; Schütt et al.,, 2017) have shown remarkable success, particularly due to their favourable linear scaling with system size and their ability to generalise across diverse chemical domains (Wood et al.,, 2025), including materials, catalysts, and molecular systems. In MPNNs-based MLIPs, atomic numbers $\bm{Z}$ are usually encoded for building node features, while positions $\bm{r}$ are used to process geometric information. Despite achieving high accuracy in predicting PES across a variety of atomic configurations, MLIPs often fail to generalise to out-of-distribution (OOD) data, yielding unreliable energy predictions when presented with atomic environments that differ significantly from those seen during training, e.g. out-of-equilibrium structures, or rare chemical compositions. However, only a few works have addressed UQ in MLIPs. Notably, Tan et al., (2023) demonstrated that Deep Ensembles remain the most effective across a variety of UQ methods, both in terms of predictive accuracy and uncertainty estimation. Despite their performance, Deep Ensembles suffer from poor efficiency, with training and inference costs scaling linearly with the number of ensemble members. Moreover, for large-scale foundation models, constructing and maintaining an ensemble is often infeasible due to the high computational and storage requirements, meaning that only a single pretrained model may be available for downstream tasks.
@@ 3.2 Variational Adaptive Dropout
where $\bm{\phi}$ denotes the parameters of the approximate distribution. Following Coscia et al., (2025), we employ an efficient parameterisation based on Variational Dropout (Kingma et al.,, 2015), which enables scalable inference in large MPNN architectures thanks to the local-reparametrization trick (Kingma et al.,, 2015) (see Algorithm 1).
Specifically, in each message-passing layer $l$, the message and update functions can be parametrised by a multi-layer perceptron (MLP), which may consist of multiple successive linear layers. For example, the message function at layer $l$ could be represented as an MLP with $S$ linear layers. We model the weights of each such internal linear layer $s=1,\dots,S$ using a Gaussian distribution [^1] We assume the same number of internal layers for both message and update networks for notational simplicity, though the approach trivially generalises. We also factorise across this dimension.:
@@
$\alpha_{l}=\alpha_{l}(\mathbf{h}^{0},a_{ij}),\quad\beta_{l}=\beta_{l}(\mathbf{h}^{0}),$ (6)
and are computed per edge and node, respectively.
This can be interpreted as injecting Gaussian noise into the weights, scaled by the corresponding variational dropout coefficients:
@@
where $\odot$ denotes element-wise multiplication. Due to the Gaussian assumption, the local reparameterization trick (Kingma et al.,, 2015) can be employed, allowing us to sample activations directly instead of weights. This drastically reduces the number of random samples per iteration, as we only need to sample once for every neuron activation rather than once for every weight, leading to both lower variance in the gradient estimates and improved computational efficiency (see Appendix D).
@@
$\begin{split}\mathcal{L}(\bm{\phi})&=\mathbb{E}_{q_{\bm{\phi}}}[\log p(\mathbf{y}\mid\mathbf{h}^{0},a_{ij},\{\bm{\omega}^{M}_{l},\bm{\omega}^{U}_{l}\}_{l=0}^{L-1})]\\ &\quad-\sum_{l=0}^{L-1}\left(D_{\mathrm{KL}}\big{[}q_{\bm{\phi}}(\bm{\omega}^{M}_{l}\mid\mathbf{h}^{0},a_{ij})\,\|\,p(\bm{\omega}^{M}_{l})\big{]}+D_{\mathrm{KL}}\big{[}q_{\bm{\phi}}(\bm{\omega}^{U}_{l}\mid\mathbf{h}^{0})\,\|\,p(\bm{\omega}^{U}_{l})\big{]}\right),\end{split}$ (8)
where the expectation is taken over all stochastic weights $\{\bm{\omega}^{M}_{l},\bm{\omega}^{U}_{l}\}_{l=0}^{L-1}$. In practice, we approximate the expectation using a single Monte Carlo sample from the variational posterior at each training iteration.
@@
$\omega^{M}_{ij;ls}$ $=\epsilon^{M}_{ij;ls}\cdot|\theta^{M}_{ij;ls}|,\quad\epsilon^{M}_{ij;ls}\sim\mathcal{N}\left(0,\,\tfrac{p}{1-p}\right),$ (9)
$\omega^{U}_{ij;ls}$ $=\epsilon^{U}_{ij;ls}\cdot|\theta^{U}_{ij;ls}|,\quad\epsilon^{U}_{ij;ls}\sim\mathcal{N}\left(0,\,\tfrac{p}{1-p}\right),$
for each message-passing layer $l$ and internal linear layer $s$. This choice yields closed-form KL divergences for both terms in eq. (8), as reported in Appendix A.1. Algorithm 2 shows how to train MLIPs using BLIP.
@@ 3.3 Equivariance and Invariance Analysis
Many MLIP architectures are explicitly designed to be either equivariant or invariant to certain input transformations (Cohen and Welling,, 2016; Weiler et al.,, 2023; Bronstein et al.,, 2017; Cohen et al.,, 2019). An equivariant network ensures that when the input is transformed (e.g., via translation or rotation), the output transforms accordingly. An invariant network is a special case of equivariant one, where it produces the same output regardless of such transformations. These properties are crucial when modelling atomic systems, where geometric symmetries should not alter the model’s predictions. BLIP can also be used in conjunction with equivariant or invariant MLIPs without violating their symmetry properties [^2] Although the model is equivariant in distribution, individual forward passes may break exact equivariance or invariance due to the injected noise. For a given input, different noise samples (which themselves do not need to be equivariant) can yield different outputs. However, when averaged across samples, the statistical moments remain equivariant.. To ensure this, the variational dropout coefficients $\alpha$ and $\beta$ must themselves be invariant, which can be achieved by implementing the inference network $E_{\bm{\psi}}$ as an invariant neural network. When $\alpha$ and $\beta$ are invariant to input transformations, the local reparameterization trick can be safely applied to any equivariant networks whose learnable layers can be expressed as unconstrained linear transformations without an irreducible representation (irrep) dimension. This includes models such as those in (Satorras et al.,, 2021; Schütt et al.,, 2021; Köhler et al.,, 2019; Schütt et al.,, 2017). However, in its current form, it excludes other important classes of equivariant MLIPs, such as MACE (Batatia et al.,, 2022) and UMA (Wood et al.,, 2025), thereby motivating future work to bridge this gap.

### v2

@@ 1 Introduction
Deep learning is transforming computational chemistry and materials science, enabling advances in molecular design (Özçelik et al., 2024; Eijkelboom et al., 2024; Zeni et al., 2023; Merchant et al., 2023), simulation-based modelling (Batatia et al., 2022; Wood et al., 2025; Neumann et al., 2024), and particle-based systems (Kipf et al., 2018; Alkin et al., 2024; Zhdanov et al., 2025). At the forefront of this progress are Message Passing Neural Networks (MPNNs) (Scarselli et al., 2008; Kipf and Welling, 2016; Gilmer et al., 2017), which have emerged as core architectures for learning in atomistic and molecular domains. Their ability to natively incorporate essential physical symmetries, such as translational, rotational, and permutation invariance or equivariance (Cohen and Welling, 2016; Weiler et al., 2023; Bronstein et al., 2017; Cohen et al., 2019), makes them particularly well-suited for modelling the complex interactions that define atomic-scale systems. Consequently, MPNNs are now widely adopted in the development of Machine Learning Interatomic Potentials (MLIPs), where they serve as efficient and accurate surrogates for quantum and classical simulations. These models enable large-scale atomistic simulations with near ab initio accuracy, dramatically accelerating research in materials discovery, catalysis, and molecular dynamics (Wood et al., 2025; Neumann et al., 2024; Batatia et al., 2022). However, like most deep learning models, MLIPs are vulnerable to failure, particularly when faced with out-of-distribution data. This issue is further worsened in scientific domains, where generating labelled data often requires expensive and time-consuming simulations (Papamarkou et al., 2024). Indeed, for simulation-based chemistry, training data is usually obtained from Density Functional Theory (DFT) calculations, which scale cubically with system size, making it impractical to generate data for larger molecules or materials, causing inevitable distributional shifts at test time (Özçelik et al., 2025; Tan et al., 2023). These challenges make principled Uncertainty Quantification (UQ) essential for enabling error-aware simulations or guiding active learning pipelines. Among existing approaches, Deep Ensembles (Lakshminarayanan et al., 2017) are commonly used for UQ in MLIPs and have shown strong empirical performance compared to other UQ methods (Tan et al., 2023). However, they require training and storing multiple models, making them computationally and memory-intensive, particularly in large-scale or high-throughput scientific workflows. Other simpler UQ methods, such as MC Dropout (Gal and Ghahramani, 2016) or mixture models, tend to be less memory-intensive but provide worse uncertainty estimates, and their predictive accuracy may degrade compared to standard deterministic models.
@@
To overcome these limitations, we introduce Bayesian Learned Interatomic Potentials (BLIPs), a general variational Bayesian method that is scalable, accurate, and provides principled, well-calibrated uncertainty estimates for MLIPs (see Figure 1). BLIP can be applied both when training and fine-tuning pretrained machine learning interatomic potentials. It builds on an adaptive version of Variational Dropout (Kingma et al., 2015), where the MPNNs’ weights are stochastically perturbed depending on the network’s input data. BLIP does not depend on the chosen MPNN architecture and seamlessly integrates with a vast family of invariant and equivariant models. Importantly, our approach adds minimal computational overhead compared to deterministic models in both training and predictive inference. We evaluate our approach on various simulation-based chemistry tasks, focusing on small and large-scale systems. In both domains, we show that training with BLIP not only enhances predictive performance but also produces meaningful uncertainty estimates. Furthermore, we demonstrate that fine-tuning a single pretrained MLIP with BLIP consistently improves predictive performance over single and ensemble models finetuning, while simultaneously providing good uncertainty estimates.
@@ 2.1 Message Passing Neural Networks
Message Passing Neural Networks (Scarselli et al., 2008; Kipf and Welling, 2016; Gilmer et al., 2017) are a class of (permutation equivariant) neural networks that process graph-structured data. Given a Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $v_{i}\in\mathcal{V}$, and edges $e_{ij}\in\mathcal{E}$, the MPNNs operations for each layer $l=0,\dots,L-1$ can be summarised as:
@@
where $\bm{h}_{i}^{l}$ is the feature vector for node $i$ at layer $l$, $a_{ij}$ are the edge attributes, and $\bigoplus_{j\in\mathcal{N}(i)}$ is a permutation invariant pooling operation over the node $i$ neighbors $\mathcal{N}(i)$. Finally, $M_{l},U_{l}$ are learnable message and update functions usually approximated with small neural networks.
@@ 2.2 Bayesian Models and Uncertainty Quantification
Bayesian models have a long history in Deep Learning (Hinton and Van Camp, 1993; Welling and Teh, 2011; Graves, 2011; Blundell et al., 2015; Gal and Ghahramani, 2016), providing a principled framework for modelling uncertainty through probabilistic inference. Instead of treating neural network weights as fixed parameters, Bayesian approaches represent them as probability distributions, $\bm{\omega}\sim p(\bm{\omega})$. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ of $N$ i.i.d. input-output observations, the goal is to infer the posterior distribution over weights, $p(\bm{\omega}\mid\mathcal{D})$, using Bayes’ theorem (Bayes, 1763). However, due to the high dimensionality of weight spaces in modern deep networks, exact Bayesian inference is computationally intractable, necessitating approximate methods. Among these, Deep Ensembles approximate the posterior by training multiple networks with different random initialisations and aggregating their predictions (Lakshminarayanan et al., 2017) and have been shown to provide state-of-the-art results in many applications (Fort et al., 2019; Tan et al., 2023). Other widely used approaches include Variational Inference (Blundell et al., 2015; Kingma et al., 2015), which approximates the posterior with a tractable parametric distribution, or Monte Carlo Dropout, which interprets dropout at test time as Bayesian inference (Gal and Ghahramani, 2016).
@@ 2.3 Machine Learning Interatomic Potentials
Density Functional Theory (DFT) (Hohenberg and Kohn, 1964; Kohn and Sham, 1965) is a widely used quantum mechanical method for calculating a range of properties for various atomic systems. Under the Born-Oppenheimer approximation (Born and Heisenberg, 1985), the potential energy surface (PES) $E$ of an atomic system depends only on positions $\bm{r}$ and atomic numbers $\bm{Z}$, with forces given by $\bm{F}=-\nabla_{\bm{r}}E(\bm{r},\bm{Z})$. However, DFT calculations are computationally expensive due to their cubic scaling with respect to the number of electrons; therefore, Machine Learning Interatomic Potentials (MLIPs) based on (equivariant) MPNNs (Batatia et al., 2022; Fu et al., 2025; Neumann et al., 2024; Rhodes et al., 2025; Schütt et al., 2017) have emerged as efficient surrogates showing remarkable success, particularly due to linear scaling with system size and ability to generalise across chemical domains (Wood et al., 2025). Despite these advances, MLIPs often fail to generalise to out-of-distribution data, such as out-of-equilibrium structures or rare chemical compositions, highlighting the need for principled UQ.
2.4 Uncertainty Quantification in Machine Learning Interatomic Potentials
Uncertainty quantification in MLIPs plays a central role in ensuring model reliability (Thaler et al., 2023; Tan et al., 2023; Chmiela et al., 2019; Kreiman and Krishnapriyan, 2025), as well as in guiding efficient exploration of chemical space through active learning (Kulichenko et al., 2024; Jinnouchi et al., 2020; Zaverkin et al., 2024). Among existing approaches, Deep Ensembles MLIPs (Tan et al., 2023; van der Oord et al., 2023; Kulichenko et al., 2023) remain one of the most effective UQ methods in terms of both predictive accuracy and uncertainty estimation. However, Deep Ensembles are computationally expensive, with training and inference costs scaling linearly with ensemble size (one model needs to be trained/queried for each ensemble member). Moreover, they are known to exhibit calibration issues (Pernot, 2022), often requiring post-hoc calibration techniques (Hu et al., 2022). Finally, for large-scale foundation models, constructing and maintaining multiple ensemble members becomes typically infeasible due to substantial computational and storage demands, leaving only a single pretrained model available for downstream tasks. For these reasons, we present a Bayesian approach that is scalable, requiring only a single trained model, and empirically exhibits good calibration, offering a practical and efficient solution for UQ in MLIPs.
@@ 3.2 Variational Adaptive Dropout
where $\bm{\phi}$ denotes the parameters of the approximate distribution. Following Coscia et al. (2025), we employ an efficient parameterisation based on Variational Dropout (Kingma et al., 2015), which enables scalable inference in large MPNN architectures thanks to the local-reparametrization trick (Kingma et al., 2015) (see Algorithm 1). Specifically, in each message-passing layer $l$, the message and update functions can be parametrised by a multi-layer perceptron (MLP), which may consist of multiple successive linear layers. For example, the message function at layer $l$ could be represented as an MLP with $S$ linear layers. We model the weights of each such internal linear layer $s=1,\dots,S$ using a Gaussian distribution [^1] We assume the same number of internal layers for both message and update networks for notational simplicity, though the approach trivially generalises. We also factorise across this dimension.:
@@
$\alpha_{l}=\alpha_{l}(\mathbf{h}^{0},a_{ij})\in\mathbb{R}^{+},\quad\beta_{l}=\beta_{l}(\mathbf{h}^{0})\in\mathbb{R}^{+},$ (6)
and are computed per edge and node, respectively by first obtaining a number $r\in[0,1]$ and then applying the transformation $\frac{r}{1-r}$. This can be interpreted as injecting Gaussian noise into the weights, scaled by the corresponding variational dropout coefficients:
@@
where $\odot$ denotes element-wise multiplication. Due to the Gaussian assumption, the local reparameterization trick (Kingma et al., 2015) can be employed, allowing us to sample activations directly instead of weights. This drastically reduces the number of random samples per iteration, as we only need to sample once for every neuron activation rather than once for every weight, leading to both lower variance in the gradient estimates and improved computational efficiency (see Appendix D).
@@
$\begin{split}\mathcal{L}(\bm{\phi})&=\mathbb{E}_{q_{\bm{\phi}}}[\log p(\mathbf{y}\mid\mathbf{h}^{0},a_{ij},\{\bm{\omega}^{M}_{l},\bm{\omega}^{U}_{l}\}_{l=0}^{L-1})]\\ &\quad-\lambda\sum_{l=0}^{L-1}\left(D_{\mathrm{KL}}\big[q_{\bm{\phi}}(\bm{\omega}^{M}_{l}\mid\mathbf{h}^{0},a_{ij})\,\|\,p(\bm{\omega}^{M}_{l})\big]+D_{\mathrm{KL}}\big[q_{\bm{\phi}}(\bm{\omega}^{U}_{l}\mid\mathbf{h}^{0})\,\|\,p(\bm{\omega}^{U}_{l})\big]\right),\end{split}$ (8)
where the expectation is taken over all stochastic weights $\{\bm{\omega}^{M}_{l},\bm{\omega}^{U}_{l}\}_{l=0}^{L-1}$, and $\lambda$ is the KL weighting, as normally done in variational inference optimisation (Higgins et al., 2017). In practice, we approximate the expectation using a single Monte Carlo sample from the variational posterior at each training iteration.
@@
$p(\bm{\omega}^{M}_{ls})=\mathcal{N}\left(\bm{0},\,\frac{p}{1-p}(\bm{\theta}^{M}_{ls})^{2}\right),\quad p(\bm{\omega}^{U}_{ls})=\mathcal{N}\left(\bm{0},\,\frac{p}{1-p}(\bm{\theta}^{U}_{ls})^{2}\right),$ (9)
for each message-passing layer $l$ and internal linear layer $s$. This choice yields closed-form KL divergences for both terms in eq. equation 8, as reported in Appendix A.1. Algorithm 2 shows how to train MLIPs using BLIP.
@@ 3.3 Equivariance and Invariance Analysis
Many MLIP architectures are explicitly designed to be either equivariant or invariant to certain input transformations (Cohen and Welling, 2016; Weiler et al., 2023; Bronstein et al., 2017; Cohen et al., 2019). An equivariant network ensures that when the input is transformed (e.g., via translation or rotation), the output transforms accordingly. An invariant network is a special case of equivariant one, where it produces the same output regardless of such transformations. These properties are crucial when modelling atomic systems, where geometric symmetries should not alter the model’s predictions. BLIP can also be used in conjunction with equivariant or invariant MLIPs without violating their symmetry properties [^2] Although the model is equivariant in distribution, individual forward passes may break exact equivariance or invariance due to the injected noise. For a given input, different noise samples (which themselves do not need to be equivariant) can yield different outputs. However, when averaged across samples, the statistical moments remain equivariant.. To ensure this, the variational dropout coefficients $\alpha$ and $\beta$ must themselves be invariant, which can be achieved by implementing the inference network $E_{\bm{\psi}}$ as an invariant neural network. When $\alpha$ and $\beta$ are invariant to input transformations, the local reparameterization trick can be safely applied to any equivariant networks whose learnable layers can be expressed as unconstrained linear transformations without an irreducible representation (irrep) dimension. This includes models such as those in (Satorras et al., 2021; Schütt et al., 2021; Köhler et al., 2019; Schütt et al., 2017). However, in its current form, it excludes other important classes of equivariant MLIPs, such as MACE (Batatia et al., 2022) and UMA (Wood et al., 2025), thereby motivating future work to bridge this gap. A more in-depth discussion is given in Appendix A.3.